OpenStudy (anonymous):

(4/9)^(2x-3) = (8/27) ^(x-1)

OpenStudy (anonymous):

please help me:((

jhonyy9 (jhonyy9):

4/9 = (2/3) on exponent 2 8/27 = (2/3) on exponent 3 so hence we get 2(2x-3)=3(x-1) 4x-6=3x-3 x=3